OpenStudy (aaronandyson):

3x - 8y = -2 9x + 4y = 8 Solve the above equations.

ganeshie8 (ganeshie8):

familiar with elimination ?

OpenStudy (aaronandyson):

i kinda forget it not practised for a long time.

OpenStudy (wolf1728):

A) 3x - 8y = -2 B) 9x + 4y = 8 Multiply equation "A" by -3 A) -9x +24y = +6 then add it to equation B B) 9x + 4Y = 8 which equals 28Y = 14 From here, solving for "X" and "Y" is easy